OpenStudy (anonymous):

Please help me! I want to know, how will understand which equation represent the relationship at any given table?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mathslover

OpenStudy (phi):

in general, this is a hard problem for a table of "random" numbers. but you could try this: plot the (x,y) points and see if they look like a curve you know. A nice "curve" would be a straight line. or maybe a circle or ellipse or parabola.

OpenStudy (phi):

if they give you a choice of equations, you could pick an (x,y) pair from the table and see if the equation is true when you substitute in the x and y numbers.

OpenStudy (anonymous):

Is there any way to check those?

OpenStudy (anonymous):

to identify?

OpenStudy (phi):

do you have a specific question that we can do ?

OpenStudy (anonymous):

um dying to post but i cant. Because of new prob with Flvs and Openstudy

OpenStudy (anonymous):

thats y, i m asking like this

OpenStudy (anonymous):

all are something like this y = - 1/2 x - 1

OpenStudy (anonymous):

i meant option were like that

OpenStudy (anonymous):

where i made error

OpenStudy (phi):

what are you given ? a table of numbers and choices for the equation ?

OpenStudy (anonymous):

in the table of number i have to choose which is the correct equation that represent in table

OpenStudy (anonymous):

they shown table, told me the select one that represent

OpenStudy (anonymous):

relation of the equation in that table

OpenStudy (phi):

there are 2 ways to do this problem if all the equations are equations of a line (in the form y= mx+b) one way is use two (x,y) pairs from the table and find the slope and y-intercept. the other way is pick an x from the table, and put that number into the first equation and see if you get the y value that is in the table. You will have to check that only one of the equations works. If 2 or more equations work, you have to test another (x,y) pair from the table, until you find the one equation that works.

OpenStudy (anonymous):

x - y -4 = -3 -2 = -2 0 = -1 2 = 0 4 = 1

OpenStudy (phi):

for example, say you had y = - 1/2 x - 1 y = 1/2 x - 1 and the table had numbers 0, -1 1, -1/2 if we try the first pair: (0,-1). that says try x=0 and see if we get -1 in the first choice: y = -1/2 * 0 -1 this gives y= -1 so this might be the equation in the 2nd equation: y= 1/2*0 -1 this is -1 also. so we have to check another point now try 1,-1/2 . first eq: y= -1/2 * 1 -1 = -1/2 - 1 = -3/2 this is not the y value in the table. so it's not this eq. try the 2nd eq.: y= 1/2 * 1 -1 = 1/2 - 1 = -1/2 this matches. so this is the answer.

OpenStudy (anonymous):

table are like this, all this number are contain that table

OpenStudy (anonymous):

thank you so much for helping@phi ! :) ur tremendously amazing!

OpenStudy (anonymous):

I love so much learning with you guys, i do not know my instructor ever let me to learn here, it is so great learning with you, i truly really enjoy and understand better here. But i do not know why they think, we are using it to cheat

OpenStudy (anonymous):

I just posted question from test which were from old tests, so that i can understand where i make error, my instructor cant help us all the time, u guys are the only friends and helper where i ask and talk. And they are stopping me to that

OpenStudy (anonymous):

I will gonna try to talk and explain them, i hope that they will understand and allow me to use Os. I am sorry @phi for talking alot. actually i m extremely mad for it. thank you so so much for helping!!!

OpenStudy (phi):

They might, but I think some people are getting A's on their tests, but don't really know the material. That is unfair to others who do not cheat and are working hard.

OpenStudy (phi):

good luck trying to work out something with them, but if your teacher won't let you come to OS, you should know it is for a good reason (to make things fair for everybody)

OpenStudy (anonymous):

yea, thanks again